# Uncertainty-enabled machine learning for emulation of regional sea-level change caused by the Antarctic Ice Sheet

### v1

@@ 7 Discussion
7.1 Comparison to related ML sea-level applications
@@
The most similar ML application to sea-level change is the “Graph Neural Network Based EmulatOR for Glacial Isostatic Adjustment” by Lin et al., (2023). This work used a similar sea-level model (equivalently called a glacial isostatic adjustment model) to generate a training dataset of RSL change from a range of ice-sheet mass change trajectories, which was used to construct a CNN. A key difference is that their training dataset used paleo ice-sheet changes over tens of thousands of years, targeting a different application. They also utilized a graph-based spherical CNN, which allows convolution operations on a spherical manifold appropriate for representing the entire Earth. That is a promising technique for whole-Earth RSL emulation, and their emulator presumably could be trained on future AIS projections (and other sea-level contributors). For quantifying uncertainty, Lin et al., (2023) utilized an ensemble learning approach, training 30 versions of their emulator. However, their 99% prediction intervals cover the true values reportedly 94.5% of the time, suggesting that actual coverage rates are less than nominal coverage rates. Additionally, such an ensemble approach has the increased computational burden of having to refit the neural network multiple times.
@@
While this study focuses on 27 specific locations, it is of interest to build a RSL emulator that produces spatial output for the entire globe, where the dimension of the output $\bm{y}$ significantly increases thereby introducing more complexity. Our preliminary work on this global spatial emulation problem has yielded varying quality of emulated outputs. Better performance could be achieved by exploring modern NN architectures, including convolutional NNs such as U-Net (Ronneberger et al.,, 2015) and the transformer (Vaswani et al.,, 2017), which could prove beneficial by effectively capturing spatial dependencies but come at the potential cost of generating a prohibitively large training data set of simulations. An additional challenge of emulating the entire globe is ensuring spherical consistency of the emulated RSL field. While the native output of the sea-level model is a structured latitude-longitude grid, treating that as two-dimensional space would ignore the fact that the longitude points are periodic and lines of latitude converge at the poles. Using a spherical representation of the output space (Lin et al.,, 2023) or emulating the spherical harmonic coefficients directly are avenues for addressing this issue.
@@
Beyond the small dimension of our output space, there are ways in which the training data could be improved. The sample of the ISMIP6 ensemble we used here was relatively small and exhibited a fairly limited range of potential AIS states. Recent single-model ice-sheet model ensembles have produced hundreds (e.g. DeConto et al.,, 2021; Berdahl et al.,, 2023; Coulon et al.,, 2024; Aschwanden and Brinkerhoff,, 2022) or thousands (e.g. Bulthuis et al.,, 2019) of model realizations, and by exploring parameter uncertainty produce a large spread of potential ice-sheet states. Utilizing more expansive training data would improve both the accuracy of the emulator as well as its applicability, but with the trade-off of increased cost of generating the RSL training data with the sea-level model. A more selective approach to choosing training data could also make the process more efficient — because many of the ensemble members had mass change close to zero, the emulators may have been focusing on behavior of little interest.
@@
Other areas for improvement include extending the RSL emulation to additional source of sea-level change (Greenland Ice Sheet, mountain glacier and icecaps, terrestrial water storage) and considering more complex Earth structure. As mentioned above, lateral variations in mantle viscosity and lithospheric thickness (Barletta et al.,, 2018; Nield et al.,, 2014) have significant impacts on RSL derived from AIS (Pan et al.,, 2021; Powell et al.,, 2022). Models accounting for 3-D viscoelastic Earth structure are significantly more computationally expensive to run (e.g., factor of $\mathcal{O}(10^{4})$, Love et al.,, 2023), which would greatly increase the value of developing fast emulators. Love et al., (2023) demonstrated potential for ML to capture the effects of lateral variations in Earth structure at a fraction of the computational cost.
@@ 7.2 Recommendations
Our results, and the others aforementioned, have shown that NNs are viable options for sea-level change emulators. Software such as PyTorch (Paszke et al.,, 2019), amongst others, has enabled the straightforward use of a variety of neural network architectures and different types of loss functions, creating many possibilities for the scientist.
Despite their substantial predictive capabilities, designing the architectures of both the NN and CVAE, including parameters like the number of layers ($L$), the quantity of units ($d_{\ell}$), the selection of activation functions, and so on, may not be straightforward and can be application specific. The process of fitting multiple models with varying architectures to pinpoint an optimal design can be somewhat ad-hoc and time consuming. Alternatively, Neural Architecture Search (NAS) has proven to be effective in automating the architecture search (Elsken et al.,, 2019). Therefore, identifying appropriate model structure for sea-level change emulators through neural architecture search (NAS) could be helpful in constructing ML-based emulators.
The simple linear regression method of performing uncertainty quantification on top of a nonlinear ML model shares elements from the weather forecasting literature (Glahn and Lowry,, 1972; Glahn et al.,, 2009; Lovegrove and Siegert,, 2023) in addition to split conformal inference (Lei et al.,, 2018). Some advantages of the approach are that that it is relatively inexpensive to implement after the initial (nonlinear) ML model has been fit and is agnostic to the particular ML model and loss function used; it is therefore recommended as an additional approach to UQ for ML predictions.
@@ 8 Conclusion
We demonstrate the efficacy of ML models, primarily the NN and CVAE, in emulating sea-level change. By employing a postprocessing strategy using simple linear regression, both the CVAE and NN exhibit accurate prediction performance and high-quality prediction intervals. The CVAE and NN models show prediction performance comparable to a baseline GP model in terms of accuracy, while the baseline GP struggles to adequately quantify uncertainty in terms of prediction interval coverage rates.
@@
We have also highlighted the use of a simple linear regression method for deriving statistically valid uncertainty quantification using (nonlinear) ML predictions. This method of calibrating uncertainties is rooted in the weather forecasting literature (Glahn and Lowry,, 1972; Glahn et al.,, 2009; Lovegrove and Siegert,, 2023) and also shares commonalities with split-conformal inference Lei et al., (2018). The method is analogous to Platt scaling (Platt et al.,, 1999), used for calibrating probabilities of discrete, as opposed to continuous, events. Our UQ approach in this work is straightforward to implement, computationally inexpensive, and has yielded well-calibrated prediction intervals or simulated sea-level change that are of a useful length.
@@
Finally, we have demonstrated an application of the NN emulator to a large sample of synthetic ice-sheet states and showed its promise for producing probabilistic projections of RSL at different locations. Further work is necessary to develop a RSL emulator applicable to the whole Earth and that accounts for important lateral variations in Earth structure. However, ML shows promise as a tool to accelerate workflows for producing sea-level projections that are fast, reliable, and with quantifiable uncertainty.
Appendix A Appendix: Using sectors in Antarctica
@@
We examine the prediction performance of ML models using the mean ice-thickness of geographic sectors across Antarctica as inputs, similar to the work of Van Katwyk et al., (2023). The model structure is the same as described in Section 3 except with using the mean thickness value for each of the 18 regions as inputs (Figure 9).
Figure 10 indicates a substantial decrease in prediction accuracy when using sectorized regions. In addition, we also observe that the empirical coverage rates for all models deviate from the nominal rates. For example, at a significance level of $\alpha=0.05$, the empirical coverage rates are 0.913 (NN), 0.913 (CVAE), 0.924 (GP), and 0.918 (RF), respectively, and the average prediction intervals are noticeably wider (e.g., 0.170 (NN), 0.169 (CVAE), 0.181 (GP), and 0.169 (RF) at $\alpha=0.05$, respectively). This suggests that more advanced dimension-reduction techniques should be employed to preserve the information in the input data.
Figure 9: A map of sectors in Antarctica from the ISMIP6 basins. Numbers represent the basin number.
Figure 10: R-squared comparison between ML model predictions and actual simulated sea-level change values using sections in Antarctica. Values are markedly lower compared to when using the full-dimensional input vector.
@@ Acknowledgments
The authors are grateful for support to MY from the 2023 Los Alamos National Laboratory Applied Machine Learning Research Fellowship. MJH, GG, SC, HKH, and TH were supported by the DOE Office of Science Early Career Research program. SC was also supported by the Laboratory Directed Research and Development program of Los Alamos National Laboratory under project number 20210952PRD3. GG was additionally supported by the Verification and Validation subprogram of the Advanced Simulation and Computing Program at the Los Alamos National Laboratory. MY and CKW are grateful for support from the University of Maine (NIH) subcontract 00082247 for work performed at the University of Missouri. This research used resources provided by the Los Alamos National Laboratory Institutional Computing Program, which is supported by the U.S. Department of Energy National Nuclear Security Administration under Contract No. 89233218CNA000001.

### v2

@@ 8 Discussion
8.1 Comparison of kernel and machine-learning emulators
The derivation of the sensitivity kernel method is based on a physics approach that makes use of reciprocity theorems as is described in Section 4. As such, because this physics-based emulator is tailored to the problem of emulating sea-level change and makes use of symmetry specific to this system on timescales where elastic effects are dominant, it is perhaps unsurprising that it performs better than the generic ML approaches compared. In contrast, the ML approaches do not have any additional information about the underlying physics besides what is represented in the training data set input-output pairs. The sensitivity kernel emulator appears to perform better than the ML emulators based on statistical tests, though the improvement in accuracy is not, for instance, several orders of magnitude. Nonetheless, the kernel method does extrapolate better for extreme input values than the ML methods. All three emulators evaluate predictions in a comparable time frame (i.e., less than 0.5 seconds), but the overhead for generating training simulations for training the ML methods is more than 20 hours whereas the kernel training just take a forward solve (i.e., a minute).
8.2 Emulator limitations and future work
@@
The most similar ML application to sea-level change is the “Graph Neural Network Based EmulatOR for Glacial Isostatic Adjustment” by lin2023georgia. This work used a similar sea-level model (equivalently called a glacial isostatic adjustment model) to generate a training dataset of RSL change from a range of ice-sheet mass change trajectories, which was used to construct a convolutional neural network (CNN). A key difference is that their training dataset used paleo ice-sheet changes over tens of thousands of years, targeting a different application. They also utilized a graph-based spherical CNN, which allows convolution operations on a spherical manifold appropriate for representing the entire Earth. That is a promising technique for whole-Earth RSL emulation, and their emulator presumably could be trained on future AIS projections (and other sea-level contributors). For quantifying uncertainty, lin2023georgia utilized an ensemble learning approach, training 30 versions of their emulator. However, their 99% prediction intervals cover the true values reportedly 94.5% of the time, suggesting that actual coverage rates are less than nominal coverage rates. Additionally, such an ensemble approach has the increased computational burden of having to refit the neural network multiple times during training.
@@
While this study focuses on 27 specific locations, it is of interest to build a RSL emulator that produces spatial output for the entire globe, where the dimension of the output $\bm{y}$ significantly increases thereby introducing more complexity. Our preliminary work on this global spatial emulation problem has yielded varying quality of emulated outputs. Better performance could be achieved by exploring modern NN architectures, including CNNs such as U-Net (ronneberger2015u) and the transformer (vaswani2017attention), which could prove beneficial by effectively capturing spatial dependencies but come at the potential cost of requiring a prohibitively large training dataset of simulations. An additional challenge of emulating the entire globe is ensuring spherical consistency of the emulated RSL field. While the native output of the sea-level model is a structured latitude-longitude grid, treating that as two-dimensional space would ignore the fact that the longitude points are periodic and lines of latitude converge at the poles. Using a spherical representation of the output space (lin2023georgia) or emulating the spherical harmonic coefficients directly are avenues for addressing this issue.
@@
Beyond the small dimension of our output space, there are ways in which the training data could be improved. The sample of the ISMIP6 ensemble we used here was relatively small and exhibited a fairly limited range of potential AIS states. Recent single-model ice-sheet model ensembles have produced hundreds (e.g. DeConto2021; Berdahl2023; Coulon2024; Aschwanden2022) or thousands (e.g. Bulthuis2019) of model realizations, and by exploring parameter uncertainty produce a large spread of potential ice-sheet states. Utilizing more expansive training data would improve both the accuracy of the emulator as well as its applicability, but with the trade-off of increased cost of generating the RSL training data with the sea-level model. A more selective approach to choosing training data could also make the process more efficient — because many of the ensemble members had mass change close to zero, the emulators may have been focusing on behavior of little interest.
@@
Other areas for improvement include extending the RSL emulation to additional source of sea-level change (Greenland Ice Sheet, mountain glacier and icecaps, terrestrial water storage) and considering more complex Earth structure. As mentioned above, lateral variations in mantle viscosity and lithospheric thickness (Barletta2018; Nield2014) have significant impacts on RSL derived from AIS (Pan2021; Powell2022). Models accounting for 3-D viscoelastic Earth structure are significantly more computationally expensive to run (e.g., factor of $\mathcal{O}(10^{4})$, Love2023), which would greatly increase the value of developing fast emulators. Love2023 demonstrated potential for ML to capture the effects of lateral variations in Earth structure at a fraction of the computational cost.
@@ 8.3 Recommendations
Our results point to the sensitivity-kernel emulator being the most accurate and requiring the least computational overhead to generate. The neural network methods we had compared against, as well as the others discussed, have shown that NNs are also viable options for sea-level change emulators. As described above, the short timescale considered here allowed us to use an elastic Earth structure for the RSL predictions. Over longer timescales where viscoelastic Earth deformation is relevant, the elastic kernel method will no longer be appropriate and the ML techniques may be more accurate. We also note than an alternative approach to the sensitivity-kernel method would be exploring forward model elastic fingerprints with reduced dimensionality of the input space. This could either be reduced spherical harmonic degree and order or representing the ice-sheet mass change as uniform within prescribed sectors. Sector-based fingerprints could be generated and subsequently scaled by mass change in each sector and then added together to generate RSL maps, exploiting the linear nature of elastic fingerprints. We did not explore this option because the sensitivity-kernel approach allowed us to retain the complete structure of the AIS thickness change field, but for applications where the global pattern of RSL is of interest, as opposed to a small number of point locations, a sector-based fingerprint approach could be advantageous.
Despite their substantial predictive capabilities, designing the architectures of both the NN and CVAE, including parameters like the number of layers ($L$), the quantity of units ($d_{\ell}$), the selection of activation functions, and so on, may not be straightforward and can be application specific. The process of fitting multiple models with varying architectures to pinpoint an optimal design can be somewhat ad-hoc and time consuming. Alternatively, NAS has proven to be effective in automating the architecture search (nas). Therefore, identifying appropriate model structure for sea-level change emulators through NAS could be helpful in constructing ML-based emulators. Moreover, software such as PyTorch (pytorch), among others, has enabled the straightforward use of a variety of neural network architectures and different types of loss functions, creating many possibilities for the scientist.
We have demonstrated the utility of two UQ methods for augmenting emulator predictions with reliable uncertainty estimates: simple linear regression and split-conformal inference. The simple linear regression method of performing UQ on top of emulator output shares elements from the weather forecasting literature (glahn1972use; glahn2009mos; lovegrove2023improving), as well as split-conformal inference (lei2018distribution). Some advantages of the approach are that that it is relatively inexpensive to implement after the initial emulator has been fit and can be used irrespective of the specific emulator and/or loss function used; it is therefore recommended as an additional approach to UQ for ML predictions in addition to split-conformal inference. In contrast, Monte Carlo dropout appears to produce overly narrow intervals for the neural network predictions in this study.
@@ 9 Conclusion
We have demonstrated the efficacy of sensitivity-kernel based methods in emulating simulated sea-level change at a substantially reduced computational cost and high accuracy. In comparison, the NN and CVAE models were close though not as accurate in emulating sea-level change. Employing UQ via split-conformal inference and a linear-regression postprocessing technique produced prediction intervals that are generally well calibrated.
@@
We have highlighted the use of both split-conformal inference and a simple linear regression method for deriving statistically valid uncertainty quantification using (nonlinear) ML predictions. The linear regression method of calibrating uncertainties is rooted in the weather forecasting literature (glahn1972use; glahn2009mos; lovegrove2023improving) and also shares commonalities with split-conformal inference (lei2018distribution). The method is analogous to Platt scaling (platt1999probabilistic), used for calibrating probabilities of discrete, as opposed to continuous, events. Both UQ approaches (linear regression postprocessing and split-conformal inference) in this work are straightforward to implement, computationally inexpensive, and have yielded well-calibrated prediction intervals of simulated sea-level change that are generally of a useful length, meaning that their scale is not too large – for instance 1.6 cm for 95 %split-conformal prediction intervals using the kernel method. Moreover, UQ is a novel contribution to the application of sensitivity kernels for SLR emulation.
@@
Finally, we have demonstrated an application of the kernel-based and NN emulators to a large sample of synthetic ice-sheet states and have shown their promise for producing probabilistic projections of RSL at different locations. Further work is necessary to develop an RSL emulator applicable to the whole Earth and that accounts for important lateral variations in Earth structure. However, the emulators constructed and compared in this work show promise as a tool to accelerate workflows for producing sea-level projections that are fast, reliable, and with quantifiable uncertainty. We expect the ML emulators to especially be valuable when longer time-scale (e.g., 2300) RSL projections are needed, where viscoelastic rheology renders the use of a kernel-based method less straightforward to implement, and the inherent nonlinearites present in that regime motivate greater use of ML models such as the neural networks presented in this work.
Figure 12: A map of sectors in Antarctica from the ISMIP6 basins where numbers represent distinct basins. This figure shows the sectors we used as an alternative approach to the flattened input vector of the neural networks. While the results of the sector-based inputs were not nearly as accurate as the flattened-vector inputs, these sectors are displayed to illustrate this alternative approach to the reader.
Appendix A Using sectors in Antarctica
@@
We examine the prediction performance of ML models using the mean ice-thickness of geographic sectors across Antarctica as inputs, similar to the work of VanKatwyk2023. The model structure is the same as described in Section 3 except with using the mean thickness value for each of the 18 regions as inputs (Figure 12). We found that the prediction accuracy, however, was substantially than using the entire flattened vector of ice thickness changes as input (i.e., $R^{2}$ less than 0.1 in all cases), and so did not pursue the sector-based approach further.
Appendix B Alternative emulators considered
Gaussian processes have been a common choice to build emulators for physical applications (higdon2008computer; gu2018robust) and hence serve as a natural comparison to NN and CVAE emulators. They are primarily used because of closed-form multivariate normal distributions conditioning on training data; see gramacy2020surrogates for exact formulations.
For the baseline GP emulator, we use a GP that maps from a reduced dimensional input space to the sea-level change at each of the 27 long-term tide gauge locations. The dimension-reduction technique we adopt in this work is discussed subsequently. The GP uses a squared-exponential kernel (williams2006gaussian; gramacy2020surrogates) whose parameters, including a nugget term (i.e., noise term), are fit with a statistical technique known as profile likelihood maximization, the mathematical steps of which are described in Chapter 5 of gramacy2020surrogates. Our implementation of profile likelihood was coded in the R programming language. While there exist some recent variants of GPs such as deep Gaussian processes (e.g., sauer2023vecchia) that are capable of handling nonstationarity, we use a stationary GP with a squared-exponential covariance function only to serve as a baseline comparison.
For baseline comparisons, we use the “RandomForestRegressor" function in a package “scikit-learn" (scikitlearn) to train an RF in Python 3.9.17. We use 100 decision trees (default) and mean squared error (MSE) as the loss function for this baseline RF. As is the case for the GP baseline, an RF is trained for each of the 27 long-term tide gauge locations using the reduced dimensional input.
While the NN and CVAE models take as input an entire vector of ice thickness changes across Antarctica and perform dimension reduction within their respective architectures, the GP and RF models use a reduced-dimension space as input. Dimension reduction was performed by projecting input vectors onto a reduced-rank principal-component basis; 250 principal components were used, which represent almost 98.7% of the total variation of the training set inputs. This technique is standard for emulating high-dimensional computer simulator output (e.g., higdon2008computer, hooten2011assessing, gopalan2022higher), but in this work we use dimension reduction on the space of inputs for the GP and RF emulators, which is likely required as the dimension of inputs grows substantially – the input vector length is on the order of 100,000 whereas there are order 1000 data vectors. For instance, computing the distance matrix needed for a GP model is an $O(n^{2}m)$ operation when there are $n$ length $m$ vectors, which for $m>>n$ can be more substantial than the $O(n^{3})$ inverse and determinant calculations needed for GPs.
Open Research
The data and code for emulators based on ML models are preserved in yoo_2025_ and developed at https://github.com/msyoostat/RSL.
The ISMIP6 ensemble dataset was retrieved from the ISMIP6 archive (https://theghub.org/dataset-listing). Scripts for preparing the ISMIP6 ensemble data for use in this study are preserved in matt_hoffman_2025_14868920 and are developed at https://github.com/SLE-E3SM/ISMIP6_processing. The 1DSeaLevelModel_FWTW sea-level model used for generating the elastic sea-level fingerprints is preserved in holly_han_2025_14868955 and developed at https://github.com/MALI-Dev/1DSeaLevelModel_FWTW. The version of SLReciprocityGJI used for generating sensitivity kernels is slightly modified from (Al-Attar_2023). The version used in this study uses modified constants to be consistent with 1DSeaLevelModel_FWTW and includes additional pre and post processing scripts for the analysis workflow in this study. The modified version used in this study is preserved in david_al_attar_2025_14868996 and developed at https://github.com/matthewhoffman/SLReciprocityGJI.
@@ Acknowledgments
The authors are grateful for support to MY from the 2023 Los Alamos National Laboratory Applied Machine Learning Research Fellowship. MJH, GG, SC, HKH, and TH were supported by the DOE Office of Science Early Career Research program. SC was also supported by the Laboratory Directed Research and Development program of Los Alamos National Laboratory under project number 20210952PRD3. GG was additionally supported by the Verification and Validation subprogram of the Advanced Simulation and Computing Program at the Los Alamos National Laboratory. MY and CKW are grateful for support from the University of Maine (NIH) subcontract 00082247 for work performed at the University of Missouri. This research used resources provided by the Los Alamos National Laboratory Institutional Computing Program, which is supported by the U.S. Department of Energy National Nuclear Security Administration under Contract No. 89233218CNA000001. Simulations were performed on machines at the National Energy Research Scientific Computing Center, a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231 using NERSC award using NERSC award ERCAP0028051.